Given the function f(x) = −2x^2 + 5x − 1, and g(x) = 5x + 3 evaluate the following:
a. f(−2)
b.f(3) + 7

c. g(x − 1)

Answer by Edwin McCravy(20056)

Given the function f(x) = -2x^2 + 5x - 1, and g(x) = 5x + 3 evaluate the following:

------------------------------------

a. f(-2)

Put -2 or (-2) in place of x in f(x) = -2x^2 + 5x - 1 

f(-2) = -2(-2)^2 + 5(-2) - 1

 Leave the left side as it is, simplify the right side:

f(-2) = -2(4) - 10 - 1
f(-2) = -8 - 10 - 1
f(-2) = -19

---------------------------------------------

b. f(3) + 7 

First find f(3) and then put + 7 after it:

Put 3 or (3) in place of x in f(x) = -2x^2 + 5x - 1 

f(3) = -2(3)^2 + 5(3) - 1

f(3) = -2(9) + 15 - 1

f(3) = -18 + 15 - 1

f(3) = -4

Now add +7 to both sides

f(3) + 7 = -4 + 7

f(3) + 7 = 3

-------------------------------------------

c. g(x - 1) 

Put x - 1 or (x - 1) in place of x in g(x) = 5x + 3  

g(x - 1) = 5(x - 1) + 3 

 Leave the left side as it is, simplify the right side:


g(x - 1) = 5x - 5 + 3

g(x - 1) = 5x - 2

Edwin